To solve a quad equation
Wouldn't it be:
2x^2-4x=-2x set = to zero
2x^2-4x+2v=0
GCF 2(X-1)(X-1)
X= +1
and NOT:
2x² - 4x = 2x
2x² - 6x = 0 {subtracted 2x from both sides}
2x(x - 3) = 0 {factored out the greatest common factor, 2x}
2x = 0 or x - 3 = 0 {set each factor equal to 0}
x = 0 or x = 3 {solved each equation for x}

Answer by stanbon(75887)

To solve a quad equation
Wouldn't it be:
2x^2-4x=-2x set = to zero
2x^2-4x+2x=0
2x^2-2x = 0
GCF = 2x
2x(x-1)
2x = 0 or x-1 = 0
X = 0 or x = 1
============================
Note: Following is a different problem because the 2x on the right
side has a positive coefficient.
============================
and NOT:
2x² - 4x = 2x
2x² - 6x = 0 {subtracted 2x from both sides}
2x(x - 3) = 0 {factored out the greatest common factor, 2x}
2x = 0 or x - 3 = 0 {set each factor equal to 0}
x = 0 or x = 3 {solved each equation for x}
---
Everything above is now correct.
=====================================
Cheers,
Stan H.